26. Let the random variable Y have pmf f(y) = 5(6/y, where y = 1, 2, 3, . . . Compute E(Y). [HINT: for Irl < 1 we have Σ zrz = r/(1-r)2]
Answer #1

We know \sum_{1}^{\infty}xr^{x}= \frac{r}{(1-r)^{2}}

From the question, r = 1/6

E(Y) = \sum_{1}^{\infty} y*f(Y)= \sum_{1}^{\infty} y*5*(\frac{1}{6})^{y} = 5*\sum_{1}^{\infty} y*(\frac{1}{6})^{y} =5* \frac{\frac{1}{6}}{(1-\frac{1}{6})^{2}} = 5*\frac{1}{6}*\frac{6^{2}}{5^{2}} = \frac{6}{5}